# Geometric foundations of thermodynamics in the quantum regime

### v1

@@ V Quantum thermodynamic fiber bundle
Let us suppose that the manifold of Gibbs states generated by the fixed observables $\{A_{1},\dots,A_{n}\}$, which is given by $\mathcal{B}=\bigl\{\rho_{\boldsymbol{\lambda}}\big|\boldsymbol{\lambda}\in U\subseteq\mathbb{R}^{n}\bigr\}\subseteq\mathcal{D}^{\circ}$, has maximal dimension $n$. Then, the correspondence $\Xi$ is a smooth map onto the $n$-dimensional submanifold $\mathcal{B}$. Additionally, if $\Xi$ is a submersion, i.e. the differential $d\Xi_{p}:T_{p}\mathcal{M}\to T_{\Xi(p)}\mathcal{B}$ is surjective for all $p\in\mathcal{M}$ we can claim the following:
@@
The conditions required for the previous definition to make sense, namely that the submanifold $\mathcal{B}$ have dimension $n$ and that the map $\Xi$ be a submersion, are not guaranteed to hold. A first immediate requirement for $\mathcal{B}$ to have maximal dimension is that the observables $\{A_{1},\dots,A_{n}\}$ be linearly independent. This ensures that the submanifold generated by the exponential of the linear combinations of such observables has the correct dimension $n$. However, dividing by the trace may entail a reduction in the dimensionality of the submanifold. For example, if the set of observables only has one element $\{A_{1}\}$ and $A_{1}$ is a multiple of the identity matrix, the set $\mathcal{B}$ contains just one element. On the contrary, if $A_{1}$ is a $n\times n$ matrix having $n$ distinct eigenvalues, $\mathcal{B}$ has dimension $n$ and $\Xi$ is a submersion. Cases where the set of observables has more than one element have to be studied separately.
@@
A quantum thermodynamic fiber bundle is a tuple $(\mathcal{M},\mathcal{B},\Xi,F)$, where $\mathcal{M}$ is the quantum thermodynamic state space (total space), with $\dim=2n+1$; $\mathcal{B}\subseteq\mathcal{D}^{\circ}$ is the base space, with $\dim=n$; $F\cong\mathbb{R}^{n+1}$ is the fiber; and $\Xi:\mathcal{M}\to\mathcal{B}$ is the smooth projection (state function). The bundle satisfies local triviality: for every $\sigma\in\mathcal{B}$, there exists a neighborhood $U\subset\mathcal{B}$ and a diffeomorphism $\phi:\Xi^{-1}(U)\to U\times F$, such that the following diagram commutes
@@
The equilibrium submanifold $\mathcal{E}\subset\mathcal{M}$ is a Legendrian section of the bundle. For $\sigma\in\mathcal{B}$, the intersection $F_{\sigma}\cap\mathcal{E}$ consists of a single element, assuming the map $\boldsymbol{\lambda}\mapsto\rho_{\boldsymbol{\lambda}}$ is injective, corresponding to the unique thermodynamic coordinates $\left(S(\boldsymbol{\lambda}),\boldsymbol{a}(\boldsymbol{\lambda}),\boldsymbol{\lambda}\right)$, where $\mu_{i}=\lambda_{i}$, $S=-\operatorname{\mathrm{tr}}(\rho\ln\rho)$, and $\boldsymbol{a}=\langle\boldsymbol{A}\rangle_{\rho}$. This point satisfies the contact constraint $\eta=0$, i.e.
@@
Points in $F_{\sigma}\setminus\mathcal{E}$ represent non-equilibrium thermodynamic configurations that still yield the same physical state $\sigma$ but violate equilibrium relations, e.g. inconsistent $\boldsymbol{a}$.
The physical interpretation is that the fiber $F_{\sigma}$ is the set of all thermodynamic labels $(S,\boldsymbol{a},\boldsymbol{\lambda})$ compatible with the same physical density matrix, $\sigma$. On $\mathcal{E}$, we have equilibrium, i.e. unique $S$, correct $\boldsymbol{a}=\langle\boldsymbol{A}\rangle_{\sigma}$, $\boldsymbol{\lambda}$ as intensive parameters. Off $\mathcal{E}$, the points represent non-equilibrium, meaning same $\sigma$, but $S\neq-\operatorname{\mathrm{tr}}(\sigma\ln\sigma)$, or $\boldsymbol{a}\neq\langle\boldsymbol{A}\rangle_{\sigma}$
@@
$\begin{split}a(\lambda)&=\langle\sigma_{z}\rangle=\tanh\lambda,\\ S(\lambda)&=\log 2-\frac{1+a(\lambda)}{2}\log(1+a(\lambda))-\frac{1-a(\lambda)}{2}\log(1-a(\lambda)).\end{split}$ (28)
@@
where $\mu:\mathbb{R}^{3}\to\mathbb{R}$ is any smooth extension of the equilibrium relation $\mu(S(\lambda),a(\lambda),\lambda)=\lambda$.
Geometrically, this means that $\mathcal{M}\cong\mathbb{R}^{3}$ is a 3D volume; $\mathcal{B}\subset\mathcal{D}^{\circ}$ is a 1D curve in the Bloch ball interior; the fiber $F_{\sigma}=\Xi^{-1}(\sigma)$ over a fixed thermal state $\sigma=\rho_{\lambda_{0}}$ is the set of all $(S,a,\lambda)$ such that $\mu(S,a,\lambda)=\lambda_{0}$, i.e. a 2D surface (codimension 1) in $\mathbb{R}^{3}$.
The equilibrium submanifold $\mathcal{E}$ is the curve $\mathcal{E}=\{(S(\lambda),a(\lambda),\lambda)\mid\lambda\in\mathbb{R}\}$, which intersects each fiber $F_{\sigma}$ at exactly one point—the unique thermodynamic coordinates consistent with thermal equilibrium. Thus, multiple points in $\mathcal{M}$ map to the same physical state $\sigma$, but only one lies on $\mathcal{E}$. Points off $\mathcal{E}$ in $F_{\sigma}$ represent non-equilibrium thermodynamic descriptions of $\sigma$: same density matrix, but incorrect entropy $S$, mismatched expectation $a$, or inconsistent intensive parameter $\lambda$. This redundancy is the hallmark of the fiber bundle structure.
@@
In the quantum thermodynamic fiber bundle, this injectivity ensures that each Gibbs state $\sigma\in\mathcal{B}$ intersects the equilibrium submanifold $\mathcal{E}$ at exactly one point. Thermal equilibrium between $A$ and $B$ therefore corresponds to both systems mapping to the same fiber $F_{\rho_{A}=\rho_{B}}$, and thus to the same unique equilibrium point $p_{\sigma}\in\mathcal{E}$. The shared coordinates $\boldsymbol{\lambda}$ define a universal intensive parameter field on $\mathcal{E}$, with level sets, e.g. fixed inverse temperature $\beta$, forming transitive equilibrium classes. The zeroth law is thereby elevated from empirical observation to geometric necessity: thermal equilibrium is transitive because equilibrium is unique.
If injectivity fails, by e.g. redundant observables $A_{2}=cA_{1}$, with $c\in\mathbb{R}$, multiple $\boldsymbol{\lambda}$ yield the same $\rho$, destroying uniqueness and allowing inconsistent temperature assignments, the zeroth law breaks. Thus, injectivity is not an additional assumption but the geometric expression of thermodynamic uniqueness: a single quantum state admits one and only one equilibrium thermodynamic description. This structural fact is what allows temperature (and all intensive parameters) to be unambiguously defined, shared across systems in contact, and transitive under thermal equilibration—the very essence of the quantum zeroth law.
@@ VI Movement along fibers
A smooth path $\gamma:[0,T]\to\mathcal{M}$ is state-preserving with respect to $\sigma\in\mathcal{B}$ if $\gamma(t)\in F_{\sigma}$, for all $t\in[0,T]$, i.e. $\mu_{i}\big(S(t),\boldsymbol{a}(t),\boldsymbol{\lambda}(t)\bigr)=c_{i}$, for $i=1,\dots,n$, where $\boldsymbol{c}\in\mathbb{R}^{n}$ satisfies $\sigma=\rho_{\boldsymbol{c}}$ via the Gibbs form.
The fiber $F_{\sigma}$ is a smooth $(n+1)$-dimensional submanifold of $\mathcal{M}$, defined by $n$ independent level set constraints. Assuming injectivity of $\boldsymbol{\lambda}\mapsto\rho_{\boldsymbol{\lambda}}$, the equilibrium point in $F_{\sigma}$ is the unique intersection $p_{\sigma}:=F_{\sigma}\cap\mathcal{E}$, where $\mu_{i}=\lambda_{i}$, $a_{i}=\operatorname{\mathrm{tr}}(A_{i}\sigma)$, and $S=-\operatorname{\mathrm{tr}}(\sigma\ln\sigma)$. A relaxation process to the equilibrium is a path $\gamma:[0,\infty)\to F_{\sigma}$ with the initial condition $\gamma(0)\in F_{\sigma}\setminus\mathcal{E}$, and limit $\lim_{t\to\infty}\gamma(t)=p_{\sigma}$.
The contact distribution $\ker\eta\subset T\mathcal{M}$ defines Legendre submanifolds of dimension $n$. While $F_{\sigma}$ is $(n+1)$-dimensional, its intersection with $\ker\eta$ gives reversible directions.
Remark 10.
Paths $\gamma(t)\in F_{\sigma}$ keep the density operator $\sigma$ fixed, so the expectation values $a_{i}(t)=\operatorname{\mathrm{tr}}(A_{i}\sigma)$ are target coordinates that may differ from the actual $\operatorname{\mathrm{tr}}(A_{i}\sigma)$ except on the equilibrium point $p_{\sigma}=F_{\sigma}\cap\mathcal{E}$. Thermodynamically, these paths in $F_{\sigma}$ enable computation of quantities like work or heat. The quantum first law can be expressed through changes in the expectation values, constrained by the structure of the fiber. For a quasistatic process, the work done is related to changes in $\boldsymbol{\lambda}(t)$, while heat is associated with $S^{\prime}(t)$. The contact distribution $\ker\eta\subset T\mathcal{M}$ selects the reversible directions inside $F_{\sigma}$. Any deviation from $\ker\eta$ generates positive entropy production, providing a geometric criterion for thermodynamic reversibility. Note that if $\gamma(t)\in\ker\eta$, the process is reversible since there is no entropy production associated to such process.

### v2

@@ V Quantum thermodynamic fiber bundle
Let us suppose that the manifold of Gibbs states generated by the fixed observables $\{A_{1},\dots,A_{n}\}$, which is given by $\mathcal{B}=\bigl\{\rho_{\boldsymbol{\lambda}}\big|\boldsymbol{\lambda}\in U\subseteq\mathbb{R}^{n}\bigr\}\subseteq\mathcal{D}^{\circ}$, has maximal dimension $n$. Then, the correspondence $\Xi$ is a smooth map onto the $n$-dimensional submanifold $\mathcal{B}$. Additionally, if $\Xi$ is a surjective submersion, i.e., the differential $d\Xi_{p}:T_{p}\mathcal{M}\to T_{\Xi(p)}\mathcal{B}$ is surjective for all $p\in\mathcal{M}$, we can claim the following:
@@
The conditions required for the previous definition to make sense, namely that the submanifold $\mathcal{B}$ has dimension $n$ and that the map $\Xi$ be a submersion, are not guaranteed to hold. A first immediate requirement for $\mathcal{B}$ to have maximal dimension is that the observables $\{A_{1},\dots,A_{n}\}$ be linearly independent. This ensures that the submanifold generated by the exponential of the linear combinations of such observables has the correct dimension $n$. However, dividing by the trace may entail a reduction in the dimensionality of the submanifold. For example, if the set of observables only has one element $\{A_{1}\}$ and $A_{1}$ is a multiple of the identity matrix, the set $\mathcal{B}$ contains just one element. On the contrary, if $A_{1}$ is an $n\times n$ matrix having $n$ different eigenvalues, $\mathcal{B}$ has dimension 1 and $\Xi$ is a submersion. Cases where the set of observables has more than one element have to be studied separately.
If all the level sets $\Xi^{-1}(\sigma)$ for $\sigma\in\mathcal{B}$ are diffeomorphic, the previous construction gives rise to a fiber bundle:
@@
A quantum thermodynamic fiber bundle is a tuple $(\mathcal{M},\mathcal{B},\Xi,F)$, where $\mathcal{M}$ is the quantum thermodynamic state space (total space), with $\dim=2n+1$; $\mathcal{B}\subseteq\mathcal{D}^{\circ}$ is the base space, with $\dim=n$; $F$ is the fiber; and $\Xi:\mathcal{M}\to\mathcal{B}$ is the smooth projection (state function). The bundle satisfies local triviality: for every $\sigma\in\mathcal{B}$, there exists a neighborhood $U\subset\mathcal{B}$ and a diffeomorphism $\phi:\Xi^{-1}(U)\to U\times F$, such that the following diagram commutes
@@
Under the conditions thus far assumed, the quantum thermodynamic fiber bundle is a trivial bundle, in the sense that it is globally diffeomorphic to a cartesian product $\mathbb{R}^{n}\times\mathbb{R}^{n+1}$. Indeed, the base space of the bundle, namely the submanifold of Gibbs states $\mathcal{B}$, is always contractible for a quantum system of finite-dimensional state space. It can be seen that $\mathcal{B}$ is homeomorphic to $\mathbb{R}^{n}$. As is well known, any fiber bundle over a contractible base space is necessarily trivial. In other words, $\mathbb{R}^{2n+1}$ is homeomorphic to $\mathbb{R}^{n}\times\Xi^{-1}(\sigma)$ for any $\sigma\in\mathcal{B}$. From topological considerations, it then follows that $\Xi^{-1}(\sigma)$ must also be contractible. In fact, $\Xi^{-1}(\sigma)$ can be argued to be diffeomorphic to $\mathbb{R}^{n+1}$.
The equilibrium submanifold $\mathcal{E}\subset\mathcal{M}$ is a Legendrian section of the bundle. For $\sigma\in\mathcal{B}$, the intersection $F_{\sigma}\cap\mathcal{E}$ consists of a single element, assuming the map $\boldsymbol{\lambda}\mapsto\rho_{\boldsymbol{\lambda}}$ is injective, corresponding to the unique thermodynamic coordinates $\left(S(\boldsymbol{\lambda}),\boldsymbol{a}(\boldsymbol{\lambda}),\boldsymbol{\lambda}\right)$, where $\mu_{i}=\lambda_{i}$, $S=-\operatorname{\mathrm{tr}}(\rho\log\rho)$, and $\boldsymbol{a}=\langle\boldsymbol{A}\rangle_{\rho}$. This point satisfies the contact constraint $\eta=0$, i.e.,
@@
Points in $F_{\sigma}\setminus\mathcal{E}$ represent non-equilibrium thermodynamic configurations that still yield the same physical state $\sigma$ but violate equilibrium relations, e.g., inconsistent $\boldsymbol{a}$.
The physical interpretation is that the fiber $F_{\sigma}$ is the set of all thermodynamic labels $(S,\boldsymbol{a},\boldsymbol{\lambda})$ compatible with the same physical density matrix, $\sigma$. On $\mathcal{E}$, we have equilibrium, i.e., unique $S$, correct $\boldsymbol{a}=\langle\boldsymbol{A}\rangle_{\sigma}$, $\boldsymbol{\lambda}$ as intensive parameters. Off $\mathcal{E}$, the points represent non-equilibrium, meaning same $\sigma$, but $S\neq-\operatorname{\mathrm{tr}}(\sigma\log\sigma)$, or $\boldsymbol{a}\neq\langle\boldsymbol{A}\rangle_{\sigma}$
@@
$\begin{split}a(\lambda)&=\langle\sigma_{z}\rangle=\tanh\lambda,\\ S(\lambda)&=\log(2\cosh\lambda)-\lambda\tanh\lambda.\end{split}$ (29)
@@
where $\mu:\mathbb{R}^{3}\to\mathbb{R}$ is any smooth extension of the equilibrium relation $\mu(S(\lambda),a(\lambda),\lambda)=\lambda$ satisfying injectivity over the $\lambda$ variable.
Geometrically, this means that $\mathcal{M}\cong\mathbb{R}^{3}$ is a 3D volume; $\mathcal{B}\subset\mathcal{D}^{\circ}$ is a 1D curve in the Bloch ball interior; the fiber $F_{\sigma}=\Xi^{-1}(\sigma)$ over a fixed thermal state $\sigma=\rho_{\lambda_{0}}$ is the set of all $(S,a,\lambda)$ such that $\mu(S,a,\lambda)=\lambda_{0}$, i.e., a 2D surface (codimension 1) in $\mathbb{R}^{3}$.
The equilibrium submanifold $\mathcal{E}$ is the curve $\mathcal{E}=\{(S(\lambda),a(\lambda),\lambda)\mid\lambda\in\mathbb{R}\}$, which intersects each fiber $F_{\sigma}$ at exactly one point—the unique thermodynamic coordinates consistent with thermal equilibrium. Thus, multiple points in $\mathcal{M}$ map to the same physical state $\sigma$, but only one lies on $\mathcal{E}$. Points off $\mathcal{E}$ in $F_{\sigma}$ represent non-equilibrium thermodynamic descriptions of $\sigma$: same density matrix, but incorrect entropy $S$, mismatched expectation $a$, or inconsistent intensive parameter $\lambda$. This redundancy is the hallmark of the fiber bundle structure. This fact is clearer in the following sections.
@@
In the quantum thermodynamic fiber bundle, this injectivity ensures that each Gibbs state $\sigma\in\mathcal{B}$ intersects the equilibrium submanifold $\mathcal{E}$ at exactly one point. Thermal equilibrium between $A$ and $B$ therefore corresponds to both systems mapping to the same fiber $F_{\rho_{A}=\rho_{B}}$, and thus to the same unique equilibrium point $p_{\sigma}\in\mathcal{E}$. The shared coordinates $\boldsymbol{\lambda}$ define a universal intensive parameter field on $\mathcal{E}$, with level sets, e.g., fixed inverse temperature $\beta$, forming transitive equilibrium classes. The zeroth law is then seen to be a consequence of the geometric structure: thermal equilibrium is transitive because equilibrium is unique.
If injectivity fails, by e.g., redundant observables $A_{2}=cA_{1}$, with $c\in\mathbb{R}$, multiple $\boldsymbol{\lambda}$ provide the same $\rho$, destroying uniqueness and allowing inconsistent temperature assignments, the zeroth law breaks. Thus, injectivity is not an additional assumption but the geometric expression of thermodynamic uniqueness: a single quantum state admits one and only one equilibrium thermodynamic description. This structural fact is what allows temperature (and all intensive parameters) to be unambiguously defined, shared across systems in contact, and transitive under thermal equilibration—the very essence of the quantum zeroth law.
Remark 9.
Under the assumption that the Gibbs map $\boldsymbol{\lambda}\mapsto\rho_{\boldsymbol{\lambda}}$ is injective (as required for the zeroth law, see Remark 8), the equilibrium submanifold $\mathcal{E}$ can be identified diffeomorphically with the manifold of Gibbs states $\mathcal{B}$. This diffeomorphism is explicitly realized by the map $\Phi:\mathcal{E}\longrightarrow\mathcal{B}$ given by $\rho_{\boldsymbol{\lambda}}\circ\pi$, where $\pi$ denotes the projection $(S,\boldsymbol{a},\boldsymbol{\lambda})\longmapsto\boldsymbol{\lambda}$. As a consequence of this identification, we may understand the base space of the quantum thermodynamic fiber bundle to be either $\mathcal{B}$ or $\mathcal{E}$, depending on what is most suitable for the issue being discussed.
By smoothly extending the functions $\mu_{i}$ off $\mathcal{E}$, the state function $\Xi$ becomes well-defined everywhere, with the property that its image is precisely the submanifold $\mathcal{B}$. In this extended construction, $\mathcal{B}$ appears as the distinguished $n$-dimensional submanifold of thermal equilibrium states inside $\mathcal{D}^{\circ}\supset\mathcal{B}$, and $\mathcal{E}\cong_{\Phi}\mathcal{B}\subset\mathcal{D}^{\circ}$. For every arbitrary state $\rho\in\mathcal{D}^{\circ}$ associate a unique effective thermal state $\Phi(p_{\rho})\in\mathcal{B}$, where $p_{\rho}\in\mathcal{E}$ is the equilibrium point sharing the same expectation values $\boldsymbol{a}(\rho)$ and entropy $S(\rho)$. This fact enables a consistent fiber-bundle description for every full-rank density operator in $\mathcal{D}^{\circ}$ through its projection onto the Gibbs submanifold $\mathcal{B}$.
Example 2.
Consider again the qubit with $\mathcal{H}\cong\mathbb{C}^{2}$ and observable $A=\sigma_{z}$. Take a non-thermal full-rank state, e.g., the diagonal density operator $\rho=\mathrm{diag}(0.9,0.1)$, with expectation value $a=\operatorname{\mathrm{tr}}(\rho\sigma_{z})=0.8$ and von Neumann entropy $S(\rho)=0.325$.
If we choose $\mu(S,a,\lambda)=\lambda$, the unique equilibrium intersection is the thermal Gibbs state with the same $a=0.8$, i.e., $\rho_{\lambda}=\mathrm{diag}((1+\tanh\lambda)/2,(1-\tanh\lambda)/2)$ where $\tanh\lambda=0.8$, so $\lambda=1.099$. The equilibrium labels are $S_{\mathrm{eq}}=0.420>S(\rho)$ (maximum entropy for fixed $a$), $a_{\mathrm{eq}}=0.8$, $\lambda_{\mathrm{eq}}=1.099$. Points off in $F_{\rho_{\lambda}}$ correspond to alternative thermodynamic descriptions of this same non-thermal $\rho$.
Physically, this non-thermal $\rho$ has excess free energy relative to its passive thermal counterpart: work can be extracted unitarily to reach the reordered passive state, then further to the ground. The fiber geometry quantifies distance to equilibrium, e.g., via pseudo-metric lengths for relaxation paths projecting to thermalization under a bath. A horizontal lift of a curve from $\rho$ to $\rho_{\lambda_{\mathrm{eq}}}$ in the base represents an optimal quasistatic driving toward the effective thermal state.
@@ VI Movement along fibers
A smooth path $\gamma:[0,T]\to\mathcal{M}$ is state-preserving with respect to $\sigma\in\mathcal{B}$ if $\gamma(t)\in F_{\sigma}$, for all $t\in[0,T]$, i.e., $\mu_{i}\big(S(t),\boldsymbol{a}(t),\boldsymbol{\lambda}(t)\bigr)=c_{i}$, for $i=1,\dots,n$, where $\boldsymbol{c}\in\mathbb{R}^{n}$ satisfies $\sigma=\rho_{\boldsymbol{c}}$ via the Gibbs form.
The fiber $F_{\sigma}$ is a smooth $(n+1)$-dimensional submanifold of $\mathcal{M}$, defined by $n$ independent level set constraints. Assuming injectivity of $\boldsymbol{\lambda}\mapsto\rho_{\boldsymbol{\lambda}}$, the equilibrium point in $F_{\sigma}$ is the unique intersection $p_{\sigma}:=F_{\sigma}\cap\mathcal{B}$, where $\mu_{i}=\lambda_{i}$, $a_{i}=\operatorname{\mathrm{tr}}(A_{i}\sigma)$, and $S=-\operatorname{\mathrm{tr}}(\sigma\log\sigma)$. A relaxation process to the equilibrium is a path $\gamma:[0,\infty)\to F_{\sigma}$ with the initial condition $\gamma(0)\in F_{\sigma}\setminus\mathcal{B}$, and limit $\lim_{t\to\infty}\gamma(t)=p_{\sigma}$. The contact distribution $\ker\eta\subset T\mathcal{M}$ defines Legendre submanifolds of dimension $n$. While $F_{\sigma}$ is $(n+1)$-dimensional, its intersection with $\ker\eta$ gives reversible directions.
Paths $\gamma(t)\in F_{\sigma}$ keep the density operator $\sigma$ fixed, so the expectation values $a_{i}(t)=\operatorname{\mathrm{tr}}(A_{i}\sigma)$ are target coordinates that may differ from the actual $\operatorname{\mathrm{tr}}(A_{i}\sigma)$ except on the equilibrium point $p_{\sigma}=F_{\sigma}\cap\mathcal{B}$. Thermodynamically, these paths in $F_{\sigma}$ enable computation of quantities like work or heat. The quantum first law can be expressed through changes in the expectation values, constrained by the structure of the fiber. For a quasistatic process, the work done is related to changes in $\boldsymbol{\lambda}(t)$, while heat is associated with $S^{\prime}(t)$. The contact distribution $\ker\eta\subset T\mathcal{M}$ selects the reversible directions inside $F_{\sigma}$. Any deviation from $\ker\eta$ generates positive entropy production, providing a geometric criterion for thermodynamic reversibility. Note that if $\gamma(t)\in\ker\eta$, the process is reversible since there is no entropy production associated with such a process.